OpenStudy (anonymous):

h(x) = x^2-6x Put that function into the form (x-p)^2-q where p and q are constants

OpenStudy (anonymous):

i got for p =3 and q = 9

OpenStudy (anonymous):

\[(x-3)^2=x^2-6x+9\] \[(x^2-3)^2- 9)

OpenStudy (anonymous):

p = 3 and q = 9

OpenStudy (anonymous):

is the inverse of that square root (x+9) +3

OpenStudy (anonymous):

(x−3)2=x2−6x+9 compating \[(x^2-3)^2- 9)\] p=3 q=9 ps: q=9 cause \[(x-3)^2=x^2-6x+9\] and we wish \[x^2-6x=0\]